# Scalable Message Passing Neural Networks: No Need for Attention in Large Graph Representation Learning

### v1

@@ Abstract
11footnotemark: 1
University of Oxford  22footnotemark: 2 Massachusetts Institute of Technology
33footnotemark: 3
McMaster University  44footnotemark: 4 Vector Institute  55footnotemark: 5 AITHYRA
@@ 1 Introduction
Traditionally, Graph Neural Networks (GNNs) [1] have primarily been applied to model functions over graphs with a relatively modest number of nodes. However, recently there has been a growing interest in exploring the application of GNNs to large-scale graph benchmarks, including datasets with up to a hundred million nodes [2]. This exploration could potentially lead to better models for industrial applications such as large-scale network analysis in social media, where there are typically millions of users, or in biology, where proteins and other macromolecules are composed of a large number of atoms. This presents a significant challenge in designing GNNs that are scalable while retaining their effectiveness.
@@
To this end, we take inspiration from the literature on Large Language Models (LLMs) and propose a simple modification to how GNN architectures are typically arranged. Our framework, Scalable Message Passing Neural Networks (SMPNNs), enables the construction of deep and scalable architectures that outperform the current state-of-the-art models for large graph benchmarks in transductive classification. More specifically, we find that following the typical construction of the Pre-Layer Normalization (Pre-LN) Transformer formulation [3] and replacing attention with standard message-passing convolution is enough to outperform the best Graph Transformers in the literature. Moreover, since our formulation does not necessarily require attention, our architecture scales better than Graph Transformers. Attention can also be easily incorporated into our framework if needed; however, in general, we find that adding attention, at least for large-scale graph transductive learning, only leads to marginal improvements in performance at the cost of being more computationally demanding.
@@
Our empirical observations, which demonstrate that SMPNNs can use many layers unlike traditional GNNs, are supported by recent theoretical studies on oversmoothing and oversharpening in graph convolutions [4] and Transformers [5]. These studies suggest the crucial role of residual connections in mitigating oversmoothing and low-frequency dominance in representations. It is worth noting, however, that the aforementioned works primarily approached this issue from a theoretical standpoint and did not scale to large graph datasets. Expanding upon previous theoretical studies on oversmoothing, we provide a universal approximation perspective. Specifically, we demonstrate that residual connections, such as those found in the Transformer block architecture and other newer blocks utilized in the LLM literature, like the Mamba block [6], are essential for preserving the universal approximation properties of downstream classifiers.
Contributions. We propose Scalable Message Passing Neural Networks (SMPNNs), a framework designed to scale traditional message-passing GNNs. Our main contributions are the following:
  • (i)
@@
    The SMPNN architecture can scale to large graphs thanks to its $\mathcal{O}(E)$ graph convolution computational complexity, outperforming state-of-the-art Graph Transformers for transductive learning without using global attention mechanisms. It also enables deep message-passing GNNs without suffering from oversmoothing, a problem that has traditionally limited these networks to shallow configurations.
  • (ii)
    We theoretically analyze the advantages of our architecture compared to traditional convolutions over graphs from a universal approximation perspective. Importantly, unlike previous works, we do not rely on the asymptotic convergence of the system to explain its behaviour.
  • (iii)
    We perform extensive experiments in large-graph transductive learning as well as on smaller datasets and demonstrate that our model consistently outperforms recently proposed Graph Transformers and other traditional scalable architectures. We also conduct ablations and additional experiments to test the importance of the different components of our architecture and support our theoretical claims.
@@ 2 Background
Graph Representation Learning and Notation. We denote a graph with $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents a set of nodes (or vertices), and $\mathcal{E}\subseteq(\mathcal{V}\times\mathcal{V})$ is a set of 2-tuples signifying edges (or links) within the graph. For any pair of nodes $v_{i}$ and $v_{j}$ within $\mathcal{G}$, their connection is encoded as $(v_{i},v_{j})\in\mathcal{E}$ if the edge originates from $v_{i}$ and terminates at $v_{j}$. The (one-hop) neighborhood of node $v_{i}$, constitutes the set of nodes sharing an edge with $v_{i}$, denoted by $\mathcal{N}(v_{i})=\{v_{j}|(v_{i},v_{j})\in\mathcal{E}\}$. To encode the structural connectivity amongst nodes in a graph of $N=|\mathcal{V}|$ nodes and $E=|\mathcal{E}|$ edges, an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ is employed. This adjacency matrix may adopt a weighted or unweighted representation. In the case of a weighted adjacency matrix, the entries $A_{ij}\in\mathbb{R}$ symbolize the strength of the connection, with $A_{ij}=0$ denoting absence of a connection if $(v_{i},v_{j})\notin\mathcal{E}$. Conversely, in an unweighted adjacency matrix, $A_{ij}=1$ signifies the presence of an edge, and $A_{ij}=0$ otherwise. The degree matrix $\mathbf{D}\in\mathbb{R}^{N\times N}$ is defined as the matrix where each entry on the diagonal is the row-sum of the adjacency matrix: $D_{ii}=\sum_{j}A_{ij}$. Based on this, we can define the graph Laplacian as $\mathbf{L}=\mathbf{D}-\mathbf{A}$. The normalized graph Laplacian, denoted as $\mathbf{L}_{\text{norm}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$, is a variation of the graph Laplacian that takes into account the degrees of the nodes. Additionally, within our context, we are interested in graphs with node features. Each node $v_{i}\in\mathcal{V}$ is accompanied by a $D$-dimensional feature vector $\mathbf{x}_{i}\in\mathbb{R}^{1\times D}$. The feature vectors for all nodes within the graph can be aggregated into a single matrix $\mathbf{X}\in\mathbb{R}^{N\times D}$ by stacking them along the row dimension.
Message Passing Neural Networks (MPNNs) are a class of Graph Neural Networks (GNNs). MPNNs operate by means of message passing, wherein nodes exchange vector-based messages to refine their representations, leveraging the graph connectivity structure as a geometric prior. Following [7], a message passing GNN layer $l$ (excluding edge and graph-level features for simplicity) over a graph $\mathcal{G}$ is defined as:
$\mathbf{x}_{i}^{(l+1)}=\phi\biggl{(}\mathbf{x}_{i}^{(l)},\bigoplus_{j\in\mathcal{N}(v_{i})}\psi(\mathbf{x}_{i}^{(l)},\mathbf{x}_{j}^{(l)})\biggr{)},$ (1)
where $\psi$ denotes a message passing function, $\bigoplus$ is some permutation-invariant aggregation operator, and $\phi$ is a readout or update function. Both $\psi$ and $\phi$ are learnable and typically implemented as MLPs. It is important to note that the update equation is local, and that at each layer each node only communicates with its one-hop neighborhood given by the adjacency matrix of the graph. More distant nodes are accessed by stacking multiple message-passing layers.
Graph Convolutional Networks (GCNs) implement a special type of message passing of the following form:
$\mathbf{X}^{(l+1)}=\zeta\Big{(}\tilde{\mathbf{A}}\mathbf{X}^{(l)}\mathbf{W}^{(l)}\Big{)},$ (2)
@@
where $\zeta$ is a non-linear activation function, $\mathbf{W}^{(l)}$ is a learnable weight matrix, and $\tilde{\mathbf{A}}$ is computed as a function of the degree matrix and the adjacency matrix. The features are aggregated row-wise:
$\Big{(}\tilde{\mathbf{A}}\mathbf{X}^{(l)}\Big{)}_{i}=c_{ii}\mathbf{x}_{i}^{(l)}+\sum_{j\in\mathcal{N}(v_{i})}c_{ij}\mathbf{x}_{j}^{(l)}\,;\,c_{ij}=\frac{1}{\sqrt{\mathrm{deg}(v_{i})\mathrm{deg}(v_{j})}}.$ (3)
@@
Transformers. The Transformer architecture [8] is characterized by its all-to-all pairwise communication between nodes via the scaled dot-product attention mechanism. Transformers can be seen as an MPNN on a fully-connected graph [9, 7], where all nodes are within the one-hop neighborhood of each other in every layer. This avoids issues in capturing long-range dependencies inherent to message passing, since they are all connected. However, Transformers discard the locality inductive bias of the input graph. Given input query, key, and value matrices, $\mathbf{Q}=\mathbf{X}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{X}\mathbf{W}_{K}$, $\mathbf{V}=\mathbf{X}\mathbf{W}_{V}\in\mathbb{R}^{N\times D}$ (where for simplicity we assume the feature dimensionality of these matrices is kept the same as the original node features, $\mathbf{X}\in\mathbb{R}^{N\times D}$), the self-attention operation in the encoder blocks computes:
$\mathbf{X}^{\prime}=\textrm{softmax}(\mathbf{Q}\mathbf{K}^{T}/\sqrt{D})\mathbf{V}.$ (4)
Attention is challenging to scale to large graphs due to its computational complexity of $\mathcal{O}(N^{2})$ in the number of nodes. Large-scale Graph Transformers typically replace this operation with linear attention to avoid GPU memory overflow.
@@
Large Graph Transformers. The emergence of large-scale graph-structured datasets such as social networks has led to increased recent interest in scaling GNNs to very large graphs with up to hundreds of millions of nodes. Large-scale graph learning settings are often transductive learning, where one is given a fixed graph with partially labeled nodes and attempts to infer the properties of the missing nodes. While Graph Transformers currently produce state-of-the-art results, their main challenge is the computational complexity of attention for a large number of nodes, which is equivalent to very long context windows in LLMs. Techniques aiming at solving this problem include architectures such as Nodeformer [10], which utilizes a kernelized Gumbel-Softmax operator, Difformer [11], which builds on recent advances in graph diffusion, and SGFormer [12], which utilizes a single linear “attention” operation without the softmax activation to avoid computational overhead when scaling to graphs with up to a hundred million nodes. Other architectures such as Exphormer [13] use sparse attention mechanisms based on expander graphs to aid scaling. Additionally, Exphormer also uses linear global attention with respect to a virtual node in parallel to the expander graph. Some of the aforementioned architectures rely on the GPS Graph Transformer framework [14], which combines attention with standard message-passing GNN layers and computes both in parallel in the form of a hybrid architecture. Other scalable architectures that are not based on Transformers, such as SGC [15] and SIGN [16], have been developed for large-scale graphs, but their performance is generally worse than that of Graph Transformers, as shown in Section 5.
@@ 3 Scalable Message Passing Neural Networks
3.1 Motivation
In the GNN literature, it has been shown that graph convolutions can exhibit asymptotic behaviors other than over-smoothing in the limit of many layers when equipped with residual connections [4]. This theoretical observation aligns with well-known best practices in the LLM community [8, 17, 18], which is highly experienced with both large-scale datasets and models. In this work, we aim to bridge the gap between the GNN and LLM literature and benefit from the cross-pollination of ideas.
Packaging Attention and Message-Passing. Attention has been much emphasized in the literature and, although it is certainly paramount, the importance of the other components around it should not be underestimated. Interestingly, in retrospect, the ubiquitous attention mechanism used in LLMs had already been proposed [19] before the Transformer architecture [20]; however, it was not until it was “packaged” into the Transformer block that we know today, that it really led to the recent breakthroughs in language modeling. The scaled-dot-product attention mechanism (and its variants) in Transformers can be seen as message-passing attentional GNNs over a complete graph [9, 7]. It is well-known among practitioners in the LLM community that this operation struggles to learn effectively without residual connections. Yet, traditional GNN architectures simply stack message-passing layers one after the other [21, 22, 23], which seems to contradict the modus operandi of large-scale model engineering.
@@ 3.2 Architecture Design and the Scalable Message Passing Block
The Transformer Block. The original Transformer architecture has proven surprisingly robust, and the only main modification it has undergone since its inception is the change in the location of the normalization, which is now applied before attention. Indeed, the advantages of applying layer normalization [24] before, as in the Pre-LN Transformer, have already been studied both theoretically [3] and empirically in the literature [25, 26, 27], and we will also follow this in our implementation. Before the Transformer, residual connections were already a common component in most deep architectures for vision and language tasks since being popularized by the ResNet model [28]. In the LLM literature, even when alternatives to attention have been proposed, such as Mamba [6], operations that perform token-wise communication (equivalent to message-passing between nodes) are always accompanied by residual connections, as well as linear projections, normalization, and sometimes gatings.
Message-Passing. We integrate a standard GCN layer into the Pre-LN Transformer style block. In particular, a single block applies the following sequence of transformations to the input matrix of node features $\mathbf{X}^{(l)}\in\mathbb{R}^{N\times D}$, where $N$ is the number of nodes and $D$ is the dimensionality of the feature vectors:
@@
Next a GCN layer (plus additional components) is used for nodewise local communication instead of the standard global self-attention:
@@
where $\tilde{\mathbf{A}}$, the degree-normalized adjacency matrix is $\tilde{\mathbf{A}}_{i,j}=1/\sqrt{\operatorname{deg}(v_{i})\operatorname{deg}(v_{j})}$ if $v_{i}$ and $v_{j}$ are neighbours and 00 otherwise. $\mathrm{SiLU}(x)=\frac{x}{1+e^{-x}}$ and it is applied elementwise. We also introduce a scaling factor $\alpha_{1}^{(l)}$ which is initialized at $10^{-6}$ for applying identity-style block initialization [29]. Importantly, the layer defined in equation 6 contains a residual connection, the importance of which will be theoretically justified in Section 4.
Pointwise-feedforward. The second part of the block is a pointwise transformation of the feature vectors preceded by another learnable normalization:
@@
Our Scalable Message Passing Neural Network block is thus of the form $\mathbf{X}\mapsto\mathbf{H}_{4}$. In Figure 1, we display the complete SMPNN architecture at different levels of granularity, which consists of stacking multiple instances of the aforementioned block. In Appendix A, we discuss how to augment SMPNNs with attention, although we find that this only leads to minor performance improvements, see Table 2 in Section 5.
@@ 3.3 Computational Complexity and Comparison with Graph Transformers
In Table 1, we compare the computational complexity of SMPNNs to that of various Graph Transformers in the literature. Our model’s graph convolution layer inherits the computational cost of GCNs, $\mathcal{O}(E)$, assuming a sparse representation of the adjacency matrix [30]. Graph Transformers such as SGFormer use both graph convolutions and linear attention, resulting in a total cost of $\mathcal{O}(N+E)$ in terms of nodewise communication operations. Note that although we are highlighting the complexity of graph convolution compared to that of linear attention, $\mathcal{O}(N)$, given our architecture also has other components that act pointwise such as linear layers (like all architectures in general) it is more accurate to think of the overall complexity of SMPNNs as being $\mathcal{O}(N+E)$. Unlike other models, we do not apply $\mathcal{O}(N^{3})$ pre-processing steps [31, 32, 33, 34, 14], which would be prohibitively expensive for the large graph datasets we are targeting. Additionally, our model does not require positional encodings, attention, augmented training loss, or edge embeddings to achieve competitive performance. We would like to highlight that the majority of Graph Transformers in Table 1 have been designed for smaller graphs and have only been demonstrated on datasets with thousands of nodes or fewer due to their quadratic complexity. Our main competitors are therefore NodeFormer [10], DIFFormer [11], and SGFormer [12], which have managed to scale to millions of nodes using less expressive or simplified versions of attention without quadratic complexity.
Table 1: Comparison of model components used by different Graph Transformers: Positional Encodings (PE), Multihead Attention (MA), Augmented Training Loss (ATL), and Edge Embeddings (EE). We also report whether the model uses All-Pair communication (typically implemented as some variant of attention), the Pre-processing and Training computational complexity, and the largest graph for which the method’s performance has been reported. *Random regular expander graph generation is needed, and graphs that fail to be near-Ramanujan are discarded and regenerated.
Model Model Components All-Pair Pre-processing Training Largest Demo PE MA AL EE GraphTransformer [31] $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ 0.2K Graphormer [32] $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ 0.3K GraphTrans [35] - $\mathcal{O}(N^{2})$ 0.3K SAT [33] $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ 0.2K EGT [34] $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ 0.5K GraphGPS [14] $\mathcal{O}(N^{3})$ $\mathcal{O}(N+E)$ 1.0K Gophormer [36] - $\mathcal{O}(Nsm^{2})$ 20K Exphormer [13] * $\mathcal{O}(N+E)$ 132K NodeFormer [10] - $\mathcal{O}(N+E)$ 2.0M DIFFormer [11] - $\mathcal{O}(N+E)$ 1.6M SGFormer [12] - $\mathcal{O}(N+E)$ 100M SMPNN - $\mathcal{O}(N+E)$ 100M

### v2

@@ 1 Introduction
Traditionally, Graph Neural Networks (GNNs) (Scarselli et al., 2009) have primarily been applied to model functions over graphs with a relatively modest number of nodes. However, recently there has been a growing interest in exploring the application of GNNs to large-scale graph benchmarks, including datasets with up to a hundred million nodes (Hu et al., 2020). This exploration could potentially lead to better models for industrial applications such as large-scale network analysis in social media, where there are typically millions of users, or in biology, where proteins and other macromolecules are composed of a large number of atoms. This presents a significant challenge in designing GNNs that are scalable while retaining their effectiveness.
@@
To this end, we take inspiration from the literature on Large Language Models (LLMs) and propose a simple modification to how GNN architectures are typically arranged. Our framework, Scalable Message Passing Neural Networks (SMPNNs), enables the construction of deep and scalable architectures that outperform the current state-of-the-art models for large graph benchmarks in transductive classification. More specifically, we find that following the typical construction of the Pre-Layer Normalization (Pre-LN) Transformer formulation (Xiong et al., 2020) and replacing attention with standard message-passing convolution is enough to outperform the best Graph Transformers in the literature. Moreover, since our formulation does not necessarily require attention, our architecture scales better than Graph Transformers. Attention can also be easily incorporated into our framework if needed; however, in general, we find that adding attention, at least for large-scale graph transductive learning, only leads to marginal improvements in performance at the cost of being more computationally demanding.
@@
Our empirical observations, which demonstrate that SMPNNs can use many layers unlike traditional GNNs, are supported by recent theoretical studies on oversmoothing and oversharpening in graph convolutions (Giovanni et al., 2023) and Transformers (Dovonon et al., 2024). These studies suggest the crucial role of residual connections in mitigating oversmoothing and low-frequency dominance in representations. It is worth noting, however, that the aforementioned works primarily approached this issue from a theoretical standpoint and did not scale to large graph datasets. Expanding upon previous theoretical studies on oversmoothing, we provide a universal approximation perspective. Specifically, we demonstrate that residual connections, such as those found in the Transformer block architecture and other newer blocks utilized in the LLM literature, like the Mamba block (Gu and Dao, 2023), are essential for preserving the universal approximation properties of downstream classifiers.
@@
Contributions. We propose SMPNNs, a framework designed to scale traditional message-passing GNNs. Our main contributions are the following: (i) The SMPNN architecture can scale to large graphs thanks to its $\mathcal{O}(E)$ graph convolution computational complexity, outperforming state-of-the-art Graph Transformers for transductive learning without using global attention mechanisms. It also enables deep message-passing GNNs without suffering from oversmoothing, a problem that has traditionally limited these networks to shallow configurations. (ii) We theoretically analyze the advantages of our architecture compared to traditional convolutions over graphs from a universal approximation perspective. Importantly, unlike previous works, we do not rely on the asymptotic convergence of the system to explain its behavior. (iii) We perform extensive experiments in large-graph transductive learning as well as on smaller datasets and demonstrate that our model consistently outperforms recently proposed Graph Transformers and other traditional scalable architectures. We also conduct ablations and additional experiments to test the importance of the different components of our architecture.
@@ 2 Background
Graph Convolutional Networks (GCNs) implement a special type of message passing of the following form: $\mathbf{X}^{(l+1)}=\zeta\Big(\tilde{\mathbf{A}}\mathbf{X}^{(l)}\mathbf{W}^{(l)}\Big),$ where $\zeta$ is an activation function, $\mathbf{W}^{(l)}$ is a learnable weight matrix, and $\tilde{\mathbf{A}}$ is computed as a function of the degree matrix and the adjacency matrix. The features are aggregated row-wise: $\Big(\tilde{\mathbf{A}}\mathbf{X}^{(l)}\Big)_{i}=c_{ii}\mathbf{x}_{i}^{(l)}+\sum_{j\in\mathcal{N}(v_{i})}c_{ij}\mathbf{x}_{j}^{(l)}\,;\,c_{ij}=\frac{1}{\sqrt{\mathrm{deg}(v_{i})\mathrm{deg}(v_{j})}}.$
@@
Transformers. The Transformer architecture (Vaswani et al., 2017) is characterized by its all-to-all pairwise communication between nodes via the scaled dot-product attention mechanism. Transformers can be seen as an MPNN on a fully-connected graph (Sáez de Ocáriz Borde, 2024; Bronstein et al., 2021), where all nodes are within the one-hop neighborhood of each other in every layer. This avoids issues in capturing long-range dependencies inherent to message passing, since they are all connected. However, Transformers discard the locality inductive bias of the input graph. Given input query, key, and value matrices, $\mathbf{Q}=\mathbf{X}\mathbf{W}_{Q}$, $\mathbf{K}=\mathbf{X}\mathbf{W}_{K}$, $\mathbf{V}=\mathbf{X}\mathbf{W}_{V}\in\mathbb{R}^{N\times D}$ (where for simplicity we assume the feature dimensionality of these matrices is kept the same as the original node features, $\mathbf{X}\in\mathbb{R}^{N\times D}$), the self-attention operation in the encoder blocks computes: $\mathbf{X}^{\prime}=\textrm{softmax}(\mathbf{Q}\mathbf{K}^{T}/\sqrt{D})\mathbf{V}.$ Attention is challenging to scale to large graphs due to its computational complexity of $\mathcal{O}(N^{2})$ in the number of nodes. Large-scale Graph Transformers typically replace this operation with linear attention to avoid GPU memory overflow.
@@
Large Graph Transformers. The emergence of large-scale graph-structured datasets such as social networks has led to increased recent interest in scaling GNNs to very large graphs with up to hundreds of millions of nodes. Large-scale graph learning tasks are often transductive learning settings, where one is given a fixed graph with partially labeled nodes and attempts to infer the properties of the missing nodes. While Graph Transformers currently produce state-of-the-art results, their main challenge is the computational complexity of attention for a large number of nodes, which is equivalent to very long context windows in LLMs. Techniques aiming at solving this problem include architectures such as Nodeformer (Wu et al., 2022), which utilizes a kernelized Gumbel-Softmax operator, DIFFormer (Wu et al., 2023a), which builds on recent advances in graph diffusion, and SGFormer (Wu et al., 2023b), which utilizes a single linear “attention” operation without the softmax activation to avoid computational overhead when scaling to graphs with up to a hundred million nodes. Other architectures such as Exphormer (Shirzad et al., 2023) use sparse attention mechanisms based on expander graphs to aid scaling. Additionally, Exphormer also uses linear global attention with respect to a virtual node in parallel to the expander graph. Some of the aforementioned architectures rely on the GPS Graph Transformer framework (Rampášek et al., 2022), which combines attention with standard message-passing GNN layers and computes both in parallel in the form of a hybrid architecture. Other scalable architectures that are not based on Transformers, such as SGC (Wu et al., 2019) and SIGN (Frasca et al., 2020), have been developed for large-scale graphs, but their performance is generally worse than that of Graph Transformers, as shown in Section 5.
Additional background on graph learning and message passing is provided in Appendix A.
@@ 3 Scalable Message Passing Neural Networks
In the GNN literature, it has been shown that graph convolutions can exhibit asymptotic behaviors other than over-smoothing in the limit of many layers when equipped with residual connections (Giovanni et al., 2023). This theoretical observation aligns with well-known best practices in the LLM community (Vaswani et al., 2017; Touvron et al., 2023; Brown et al., 2020), which is highly experienced with both large-scale datasets and models. In this work, we aim to bridge the gap between the GNN and LLM literature and benefit from the cross-pollination of ideas.
Packaging Attention and Message-Passing. Attention has been much emphasized in the literature and, although it is certainly paramount, the importance of the other components around it should not be underestimated. Interestingly, in retrospect, the ubiquitous attention mechanism used in LLMs had already been proposed (Bahdanau et al., 2016) before the Transformer architecture (Vaswani et al., 2017); however, it was not until it was “packaged” into the Transformer block as we know it today, that it really led to the recent breakthroughs in language modeling. The scaled-dot-product attention mechanism (and its variants) in Transformers can be seen as message-passing attentional GNNs over a complete graph (Sáez de Ocáriz Borde, 2024; Bronstein et al., 2021). It is well-known among practitioners in the LLM community that this operation struggles to learn effectively without residual connections. Yet, traditional GNN architectures simply stack message-passing layers one after the other (Kipf and Welling, 2017; Schlichtkrull et al., 2017; Veličković et al., 2018), which seems to contradict the modus operandi of large-scale model engineering.
@@ 3.1 Architecture Design and the Scalable Message Passing Block
The Transformer Block. The Transformer architecture has proven surprisingly robust, and the only main modification it has undergone since its inception is the change in the location of normalization, which is now applied before attention. The advantages of applying layer normalization (Ba et al., 2016) before, as in the Pre-LN Transformer, have already been studied both theoretically (Xiong et al., 2020) and empirically in the literature (Baevski and Auli, 2019; Child et al., 2019; Wang et al., 2019), and we will also follow this in our implementation. Before the Transformer, residual connections were already a common component in most deep architectures for vision and language tasks since being popularized by the ResNet model (He et al., 2015). In the LLM literature, even when alternatives to attention have been proposed, such as Mamba (Gu and Dao, 2023), operations that perform token-wise communication (equivalent to message-passing between nodes) are always accompanied by residual connections, as well as linear projections, normalization, and sometimes gatings.
Message-Passing. We integrate a standard GCN layer into the Pre-LN Transformer-style block. In particular, a single block applies the following sequence of transformations to the input matrix of node features $\mathbf{X}^{(l)}\in\mathbb{R}^{N\times D}$, where $N$ is the number of nodes and $D$ is the dimensionality of the feature vectors:
@@
Next, a GCN layer (plus additional components) is used for nodewise local communication instead of the standard global self-attention:
@@
where $\tilde{\mathbf{A}}$, the degree-normalized adjacency matrix, is $\tilde{\mathbf{A}}_{i,j}=1/\sqrt{\operatorname{deg}(v_{i})\operatorname{deg}(v_{j})}$ if $v_{i}$ and $v_{j}$ are neighbors and 0 otherwise. $\mathrm{SiLU}(x)=\frac{x}{1+e^{-x}}$ and it is applied elementwise. We also introduce a scaling factor $\alpha_{1}^{(l)}$ which is initialized at $10^{-6}$ for applying identity-style block initialization (Peebles and Xie, 2023). Importantly, the layer defined in equation 2 contains a residual connection, the importance of which will be theoretically justified in Section 4.
Pointwise feedforward. The second part of the block is a pointwise transformation of the feature vectors preceded by another learnable normalization:
@@
Our Scalable Message Passing Neural Network block is thus of the form $\mathbf{X}\mapsto\mathbf{H}_{4}$. In Figure 1, we display the complete SMPNN architecture at different levels of granularity, which consists of stacking multiple instances of the aforementioned block. In Appendix B, we discuss how to augment SMPNNs with attention, although we find that this only leads to minor performance improvements, see Table 2 in Section 5.
@@ 3.2 Computational Complexity and Comparison with Graph Transformers
In Table 1, we compare the computational complexity of SMPNNs to that of various Graph Transformers in the literature. Our model’s graph convolution layer inherits the computational cost of GCNs, $\mathcal{O}(E)$, assuming a sparse representation of the adjacency matrix (Kipf and Welling, 2017). Graph Transformers such as SGFormer use both graph convolutions and linear attention, resulting in a total cost of $\mathcal{O}(N+E)$ in terms of nodewise communication operations. Note that although we are highlighting the complexity of graph convolution compared to that of linear attention, $\mathcal{O}(N)$, given our architecture also has other components that act pointwise such as linear layers (like all architectures in general) it is more accurate to think of the overall complexity of SMPNNs as being $\mathcal{O}(N+E)$. Unlike other models, we do not apply $\mathcal{O}(N^{3})$ pre-processing steps (Dwivedi and Bresson, 2021; Ying et al., 2021; Chen et al., 2022; Hussain et al., 2022; Rampášek et al., 2022), which would be prohibitively expensive for the large graph datasets we are targeting. Additionally, our model does not require positional encodings, attention, augmented training loss, or edge embeddings to achieve competitive performance. We would like to highlight that the majority of Graph Transformers in Table 1 have been designed for smaller graphs and have only been demonstrated on datasets with thousands of nodes or fewer due to their quadratic complexity. Our main competitors are therefore NodeFormer (Wu et al., 2022), DIFFormer (Wu et al., 2023a), and SGFormer (Wu et al., 2023b), which have managed to scale to millions of nodes using less expressive or simplified versions of attention without quadratic complexity.
Table 1: Comparison of model components used by different Graph Transformers: Positional Encodings (PE), Multihead Attention (MA), Augmented Training Loss (ATL), and Edge Embeddings ($E$). We also report whether the model uses All-Pair communication (typically implemented as some variant of attention), the Pre-processing and Training computational complexity, and the largest graph for which the method’s performance has been reported. *Random regular expander graph generation is needed, and graphs that fail to be near-Ramanujan are discarded and regenerated.
Model Model Components All-Pair Pre-processing Training Largest Demo PE MA ATL EE GraphTransformer (Dwivedi and Bresson, 2021) $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ 0.2K Graphormer (Ying et al., 2021) $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ 0.3K GraphTrans (Jain et al., 2021) - $\mathcal{O}(N^{2})$ 0.3K SAT (Chen et al., 2022) $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ 0.2K EGT (Hussain et al., 2022) $\mathcal{O}(N^{3})$ $\mathcal{O}(N^{2})$ 0.5K GraphGPS (Rampášek et al., 2022) $\mathcal{O}(N^{3})$ $\mathcal{O}(N+E)$ 1.0K Gophormer (Zhao et al., 2021) - $\mathcal{O}(Nsm^{2})$ 20K Exphormer (Shirzad et al., 2023) * $\mathcal{O}(N+E)$ 132K NodeFormer (Wu et al., 2022) - $\mathcal{O}(N+E)$ 2.0M DIFFormer (Wu et al., 2023a) - $\mathcal{O}(N+E)$ 1.6M SGFormer (Wu et al., 2023b) - $\mathcal{O}(N+E)$ 100M SMPNN - $\mathcal{O}(N+E)$ 100M